# Scalable Multitask Learning Using Gradient-based Estimation of Task Affinity

### v1

@@ Abstract.
Multitask learning; Task Affinity Estimation; Task Grouping
@@ 1. Introduction
A key concept underlying many multitask learning algorithms is a notion of task affinity, which can capture the abovementioned positive or negative transfer effects across tasks in a precise way. For instance, one can compare pairwise task affinity (standley2020tasks; fifty2021efficiently)—the loss of a model trained on each pair of tasks—against the loss of a model trained on each task. Given a notion of task affinity, a common recipe for designing multitask learning algorithms involves (1) Task affinity computation that builds a task affinity matrix, then (2) task grouping that uses this task affinity matrix to group tasks with positive transfers together, and finally (3) multitask training that fits a separate model per task group.
The performance improvement achieved through this paradigm depends on the notion of task affinity and the grouping procedure. Moreover, the ability to leverage this paradigm hinges on the computation of task affinity (Step 1 above), which becomes expensive as the number of tasks grows. As a case in point, the computational complexity of pairwise task affinity scales quadratically with the number of tasks: this implies that even for community detection with 100 labelings, using pairwise task affinity requires training nearly 5000 models for computing the affinity matrix.
In this paper, we scale up this multitask learning paradigm by dramatically speeding up the first step of task affinity computation for two canonical examples of task affinities: Pairwise and higher-order task affinity (See Examples 2.2, 2.3). In our experiments on various real-world datasets representing different applications, our algorithm can reduce the task affinity computation time by nearly 32$\times$ compared to full model training while incurring less than $2.7\%$ error. In addition to this dramatic efficiency improvement, we also design a more robust method for task grouping (Step 2). Taken together, these new techniques match or improve the performance of previous multitask models.
@@
The primary challenge for task affinity computation is how to avoid training a large number of multitask models on various task combinations. The key technical insight behind our algorithm is to leverage a linearization property of deep neural networks, including large language models. The linearization property for a neural network means that we can approximate the model loss for a pre-trained meta-initialization and an input/output pair by using a gradient-based Taylor’s expansion centered at the meta-initialization. This linearization property has been observed for large language model fine-tuning in recent works, albeit not for the purpose of multitask learning (malladi2023kernel; wei2023ntk; malladi2023fine). Here, we leverage linearization to estimate task affinities in an efficient manner by using the first-order Taylor expansion from a pre-trained model, thereby saving the computation of backpropagation during model fine-tuning. This algorithm, Grad-TAE, is illustrated in Figure 1.
@@
In more detail, we first compute the gradient at the initialization and then map the gradients to task labels with logistic regression. The dimension of this regression can be high, especially for heavily parameterized models. Thus, we use a dimension reduction technique and apply the Johnson-Lindenstrauss Lemma to give an error analysis. On experiments of datasets with 100 tasks, we show that this approach estimates pairwise task affinity with 45$\times$ fewer FLOPs and 11$\times$ less GPU hours than fully computing the true scores, with only 5.7% relative error. For higher-order task affinity, our approach uses 32$\times$ fewer FLOPs and 5$\times$ less GPU hours, with only 2.7% relative error. Furthermore, our approach also scales to a large-scale graph with over 21M edges and 500 tasks and estimates the task affinities within 5% relative errors with 112.3 GPU hours, while computing the true affinity scores can take over 8000 GPU hours. Our algorithm is also suitable for accelerating task selection methods that are typically computationally expensive. An example is forward or backward subset selection (hastie2009elements), which is a popular heuristic but requires evaluating quadratically many task combinations.
@@
As for the second step, we design a new clustering algorithm that uses these estimated task influences efficiently through a Semi-Definite Programming (SDP) relaxation formulation. The clustering algorithm takes the estimated task affinity matrix $T$ (of size $n\times n$) & the number $k$ of task groups as input, then solves an SDP for maximizing the average density of the $k$ groups. Since the SDP is a convex program, it can be solved efficiently, and we round the resulting solution to get the final task groups. Our experiments indicate that our clustering algorithm is more robust and performant than commonly used clustering techniques such as spectral clustering (ng2001spectral) and Lloyd’s algorithm (lloyd1982least). Once we have the task groups from the clustering, we can partition the tasks into subsets and train a separate model on tasks within each subset — this overall algorithm is called Grad-TAG. In experiments, we show that our approach achieves the Pareto optimum in terms of error rate and computation cost. For multi-label prediction on graphs trained with a $3$-layer GNN, Grad-TAG achieves comparable performance with over four baselines, while using 32$\times$ fewer FLOPs and 5$\times$ less GPU hours. For instruction fine-tuning of language models using T5-Base, Grad-TAG uses 48$\times$ fewer FLOPs and 11$\times$ less GPU hours with comparable performance to the best baseline. The code repository for reproducing our experiments can be found at: https://github.com/VirtuosoResearch/ScalableMTL.
@@
Organization: We briefly touch on related work and then provide the technical preliminaries for the rest of the paper. In section 3, we outline our task affinity estimation procedure Grad-TAE, along with a theoretical error analysis for the estimation error. Then, we present the clustering approach for task grouping and the overall algorithm Grad-TAG in Section 4. Finally, we provide a thorough empirical evaluation of the Grad-TAG algorithm for a variety of multitask learning settings in Section 5.
@@ 1.1. Related Work
Multitask learning is a fundamental problem with many applications, such as federated learning (smith2017federated), road safety modeling (nippani2024graph), and language model fine-tuning (longpre2023flan). This problem has been studied since the early literature of data mining (ben2002theoretical). As the number of tasks increases, modeling task relationships becomes increasingly complex and challenging (ma2018modeling; zhu2021transfer). These relationships are influenced by data distribution characteristics, including covariate and label shifts (wu2019understanding). Thus, designing optimization algorithms for multitask learning is challenging (li2023identification; li2023boosting). We contribute to this literature by proposing a new approach to significantly speed up the computation of task affinity scores for modeling task relationships. We now proceed to discuss several lines of work most related to ours.
@@
Task Similarity Measures. Previous works (standley2020tasks; fifty2021efficiently) estimate task affinities between every pair of tasks. The computation complexity of such methods scales quadratically with the number of tasks. Another approach is to use task embeddings (vu2021spot), i.e., training one model on each task and measuring the cosine similarity between the model weights. Although this approach scales linearly with the number of tasks, the measures tend to be noisy. Intuitively, if two tasks are similar, their gradients should exhibit higher cosine similarity. This idea can be implemented to balance training by dynamically tuning gradient magnitudes (chen2018gradnorm), or to project the gradients noto the span of other tasks’ gradients that have a conflicting gradient (fifty2021efficiently). The same idea can also be implemented to choose auxiliary tasks that are most beneficial for a primary task (dery2021auxiliary). Similarity measures based on feature representations of tasks have also been applied to grouping tasks (sherif2023stg) and used to predict task transferabilities (ayman2023task). The main advantage of these approaches is their efficiency, as only a single multitask model needs to be trained. The downside is that the gradients can be noisy during a stochastic training procedure. For example, azorin2023s empirically observed that representation and gradient similarity measures do not consistently correlate with actual MTL performance. Thus, a more accurate approach is to build measures that approximate multitask outcomes directly; see recent work on designing surrogate models for multitask learning systems (li2023identification; li2023boosting).
@@
Transferability Estimation. There have also been developments on information theoretic measures of transferability in recent literature. One natural idea is to evaluate conditional entropy between target pseudo labels (assigned by a pretrained source model) and real target label (bao2019information). Log Expected Empirical Predictor (nguyen2020leep) proposes a modified procedure by using soft predictions from the source model. These methods do not utilize feature embeddings in the measure (vu2020exploring); TransRate (huang2022frustratingly) introduces a surrogate measure based on mutual information that also incorporates feature embeddings. An improved estimation method with better robustness can be achieved by shrinkage (ibrahim2022newer). In the fine-tuning setting, the distance between the model search and the pretrained initialization can indicate the level of generalization capability (li2021improved). The geometry relates to the Hessian of the loss, which has been shown to correlate with the generalization performance of fine-tuned models (ju2022robust). ju2023generalization extend this Hessian measure to graph neural networks, which can guide the design of optimization algorithms to regularize the Hessian of neural networks (ju2023noise).
@@
Multitask Learning Optimization Algorithms. Multitask learning can be viewed as a multiobjective optimization problem (papadimitriou2000approximability), where the goal is to identify the Pareto frontier among multiple objectives (sener2018multi). One common MTL optimization algorithm is to reweight task losses and optimize a weighted combination of task losses (liu2022auto; royer2024scalarization). Our goal is to maximize the averaged prediction performance of all tasks. Thus, we are interested in partitioning the tasks into similar groups, where tasks are closely related within each group and can differ significantly across groups. Another interesting line of work is designing branching neural networks such as tree structures (vandenhende2019branched; guo2020learning), where each layer contains multiple modules to handle different tasks (lu2017fully). Compared with branching methods, task grouping may be more suitable for handling a large number of tasks (like hundreds to thousands). In this regime, negative interference between tasks is almost unavoidable, and clustering tasks into similar groups could provide a more efficient strategy than designing a single neural network that handles all tasks.
@@
Influence Functions. There is a line of work on estimating the influence of adding or removing one sample on the whole dataset. Influence functions (koh2017understanding) based on efficient approximation of the Hessian inverse provide one way to approximate this. Random sampling-based approaches to measuring leave-one-out influence have also been studied (ilyas2022datamodels; park2023trak). The distinction between these works and us is we focus on task-level affinity, whereas this literature focuses on estimating the influence of a single data sample.
@@ 2. Preliminaries
Suppose we are interested in making predictions on $n$ tasks. We are given a set of samples for training and testing of each task. Our goal is to design a prediction algorithm to maximize the averaged testing performance over all the $n$ tasks simultaneously. We assume that the samples from all the tasks are supported on a joint product between a $p$-dimensional feature space $\mathcal{X}$ and a label space $\mathcal{Y}$. In order to precisely discuss task relationships, we formally define what we mean by a multitask learning algorithm.
@@
For any subset $S\subseteq\{1,2,\dots,n\}$, a multitask learning algorithm $f$ takes the training data of all the tasks in $S$ and combines them in a joint training procedure. Then, the (jointly trained) model is tested on each task $t\in S$. In the end, a test result is obtained for each $t$. Let us denote the test result as $f(S,t)$. Thus, the output of the algorithm will include a total of $\left\lvert S\right\rvert$ results for any subset $S$, one for each $t\in S$.
Given a multitask learning algorithm, the transfer between the $n$ tasks can then be viewed through the results of $f$, applied to combinations of tasks as subsets. This notion of transfer underlies many existing multitask learning systems. We give two examples below, which have been used in prior works to tackle task transfer in complex visual systems (zamir2018taskonomy; standley2020tasks).
@@
Next, we discuss higher-order task affinity, which is analogous to sampling features in random forests. First, fix an integer $m$, which is the number of subsets we would like to sample (e.g., analogous to the number of decision trees in a random forest). We independently sample $m$ subsets out of the set $\{1,2,\ldots,n\}$, each subset having a size of $\alpha$, chosen uniformly over all such subsets. Let us denoted the $m$ subsets as $S_{1},S_{2},\ldots,S_{m}$. Then, compute $f(S_{k},j)$, for every $k=1,2,\dots,m$, and $j=1,\dots,\alpha$. Lastly, compute $T_{i,j}$ as the average value of $f$ among all subsets including tasks $i,j$:
@@
where $n_{i,j}$ is the number of subsets that include both $i,j$. This leads to another task affinity matrix $T$, better capturing the higher-order relationship among tasks.
@@ 3. Task Affinity Estimation
We now describe a new method for estimating task affinity scores. To circumvent the cost of full-model training, we start by describing an empirical observation regarding pre-training and fine-tuning. Then, we present our approach to estimating fine-tuned model parameters for task subsets. Additionally, we use random projection to reduce the dimension of the gradients. We provide an error analysis to justify the design of our algorithm.
@@ 3.1. Linearization of Fine-tuned Models
Our method is motivated by the fact that once we pre-train all the $n$ tasks to obtain a meta-initialization, this initialization can provide representations that can be quickly adapted to the remaining tasks. This is based on the premise that the underlying tasks share structural similarities in multitask learning. As the model fine-tuned to a subset of tasks stays in the affinity of the initiation, the fine-tuning procedure behaves like linear models locally.
To illustrate this observation, we consider three distinct scenarios involving graph neural networks (GNNs) and transformers (BERT and T5). We test GNNs on a multi-label prediction dataset on a YouTube graph (yang2012defining), using a 3-layer SIGN network (frasca2020sign). This dataset includes $n=100$ subtasks, one corresponding to the node labels of a subgraph of the whole graph. For transformers, we take a pretrained BERT model and fine-tune it on a sentence classification dataset (yu2020fine), which contains $n=26$ tasks. We also use a pretrained T5-Base model and fine-tune it on a sentence classification dataset with $100$ instructions (bach2022promptsource), which has $n=100$ tasks. In each experiment, we first train a meta-initialization $\theta^{\star}$ by training on all tasks combined. Then, we fine-tune $\theta^{\star}$ on a random subset of the tasks.

### v2

@@ Abstract.
Multitask learning; Task Affinity Estimation; Task Grouping; Efficiency
@@ 1. Introduction
A key concept underlying many multitask learning algorithms is a notion of task affinity, which can capture the abovementioned positive or negative transfer effects across tasks precisely. For instance, one can compare pairwise task affinity (standley2020tasks; fifty2021efficiently)—the loss of a model trained on each pair of tasks—against the loss of a model trained on each task. Given a notion of task affinity, a common recipe for designing multitask learning algorithms involves (1) Task affinity computation that builds a task affinity matrix, then (2) task grouping that uses this task affinity matrix to group tasks with positive transfers together, and finally (3) multitask training that fits a separate model per task group.
The performance improvement achieved through this paradigm depends on the task affinity and the grouping procedure. Moreover, the ability to leverage this paradigm hinges on the computation of task affinity (Step 1 above), which becomes expensive as the number of tasks grows. As a case in point, the computational complexity of pairwise task affinity scales quadratically with the number of tasks: this implies that even for community detection with 100 labelings, using pairwise task affinity requires training nearly 5000 models for computing the affinity matrix.
In this paper, we scale up this multitask learning paradigm by dramatically speeding up the first step of task affinity computation for two canonical examples of task affinities: Pairwise and higher-order task affinity (See Examples 2.2, 2.3). In our experiments on various real-world datasets representing different applications, our algorithm can reduce the task affinity computation time by nearly 32$\times$ compared to full model training while incurring less than $2.7\%$ error. In addition to this dramatic efficiency improvement, we also design a more robust method for task grouping (Step 2). Together, these new techniques match or improve the performance of previous multitask models.
@@
The primary challenge for task affinity computation is avoiding training many multitask models on various task combinations. The key technical insight behind our algorithm is to leverage a linearization property of deep neural networks, including large language models. The linearization property for a neural network means we can approximate the model loss for a pre-trained meta-initialization and an input/output pair by using a gradient-based Taylor’s expansion centered at the meta-initialization. This linearization property has been observed for large language model fine-tuning in recent works, albeit not for the purpose of multitask learning (malladi2023kernel; wei2023ntk; malladi2023fine). Here, we leverage linearization to estimate task affinities in an efficient manner by using the first-order Taylor expansion from a pre-trained model, thereby saving the computation of backpropagation during model fine-tuning. This algorithm, Grad-TAE, is illustrated in Figure 1.
@@
In more detail, we first compute the gradient at the initialization and then map the gradients to task labels with logistic regression. The dimension of this regression can be high, especially for heavily parameterized models. Thus, we use a dimension reduction technique and apply the Johnson-Lindenstrauss Lemma to give an error analysis. On experiments of datasets with 100 tasks, we show that this approach estimates pairwise task affinity with 45$\times$ fewer FLOPs and 11$\times$ less GPU hours than fully computing the true scores, with only 5.7% relative error. For higher-order task affinity, our approach uses 32$\times$ fewer FLOPs and 5$\times$ less GPU hours, with only 2.7% relative error. Furthermore, our approach also scales to a large-scale graph with over 21M edges and 500 tasks. It estimates the task affinities within 5% relative errors with 112.3 GPU hours, while computing the true affinity scores can take over 8000 GPU hours. Our algorithm is also suitable for accelerating task selection methods that are typically computationally expensive. An example is forward or backward subset selection (hastie2009elements), which is a popular heuristic but requires evaluating quadratically many task combinations.
@@
As for the second step, we design a new clustering algorithm that uses these estimated task influences through a semi-definite programming (SDP) relaxation formulation. The clustering algorithm takes the estimated task affinity matrix $T$ (of size $n\times n$) & the number $k$ of task groups as input, then solves an SDP for maximizing the average density of the $k$ groups. Since the SDP is a convex program, it can be solved efficiently, and we round the resulting solution to get the final task groups. Our experiments indicate that our clustering algorithm is more robust and performant than commonly used clustering techniques such as spectral clustering (ng2001spectral) and Lloyd’s algorithm (lloyd1982least). Once we have the task groups from the clustering, we can partition the tasks into subsets and train a separate model on tasks within each subset — this overall algorithm is called Grad-TAG. Experiments show that our approach achieves the Pareto optimum regarding error rate and computation cost. For multi-label prediction on graphs trained with a $3$-layer GNN, Grad-TAG achieves comparable performance with over four baselines while using 32$\times$ fewer FLOPs and 5$\times$ less GPU hours. For instruction tuning of language models using T5-Base, Grad-TAG uses 48$\times$ fewer FLOPs and 11$\times$ less GPU hours with comparable performance to the best baseline. The code repository for reproducing our experiments can be found at https://github.com/VirtuosoResearch/ScalableMTL.
@@
Organization: We briefly touch on related work and then provide the technical preliminaries for the rest of the paper. In section 3, we outline our task affinity estimation procedure Grad-TAE, along with a theoretical error analysis for the estimation error. Then, we present the clustering approach for task grouping and the overall algorithm Grad-TAG in Section 4. Finally, we provide a thorough empirical evaluation of the Grad-TAG algorithm for various multitask learning settings in Section 5.
@@ 1.1. Related Work
Multitask learning is a fundamental problem with many applications, such as federated learning (smith2017federated), road safety modeling (nippani2024graph), and language model fine-tuning (longpre2023flan). This problem has been studied since the early literature of data mining (ben2002theoretical). As the number of tasks increases, modeling task relationships becomes increasingly complex and challenging (ma2018modeling; zhu2021transfer). These relationships are influenced by data distribution characteristics, including covariate and label shifts (wu2019understanding). Thus, designing optimization algorithms for multitask learning is challenging (li2023identification; li2023boosting). We contribute to this literature by proposing a new approach to significantly speed up the computation of task affinity scores for modeling task relationships. We now proceed to discuss several lines of work that are most related to ours.
@@
Task Similarity Measures. Previous works (standley2020tasks; fifty2021efficiently) estimate task affinities between every pair of tasks. The computation complexity of such methods scales quadratically with the number of tasks. Another approach is to use task embeddings (vu2021spot), i.e., training one model on each task and measuring the cosine similarity between the model weights. Although this approach scales linearly with the number of tasks, the measures tend to be noisy. Intuitively, if two tasks are similar, their gradients should exhibit higher cosine similarity. This idea can be implemented to balance training by dynamically tuning gradient magnitudes (chen2018gradnorm), or to project the gradients noto the span of other tasks’ gradients that have a conflicting gradient (fifty2021efficiently). The same idea can also be implemented to choose auxiliary tasks most beneficial for a primary task (dery2021auxiliary). Similarity measures based on feature representations of tasks have also been applied to grouping tasks (sherif2023stg) and used to predict task transferabilities (ayman2023task). The main advantage of these approaches is their efficiency, as only a single multitask model needs to be trained. The downside is that the gradients can be noisy during a stochastic training procedure. For example, azorin2023s empirically observed that representation and gradient similarity measures do not consistently correlate with actual MTL performance. Thus, a more accurate approach is to build measures that approximate multitask outcomes directly; see recent work on designing surrogate models for multitask learning systems (li2023identification; li2023boosting).
@@
Transferability Estimation. There have also been developments on information theoretic measures of transferability in recent literature. One natural idea is to evaluate conditional entropy between target pseudo labels (assigned by a pretrained source model) and real target label (bao2019information). Log Expected Empirical Predictor (nguyen2020leep) proposes a modified procedure using soft predictions from the source model. These methods do not utilize feature embeddings in the measure (vu2020exploring); TransRate (huang2022frustratingly) introduces a surrogate measure based on mutual information that also incorporates feature embeddings. An improved estimation method with better robustness can be achieved by shrinkage (ibrahim2022newer). In the fine-tuning setting, the distance between the model search and the pretrained initialization can indicate the level of generalization capability (li2021improved). The geometry relates to the Hessian of the loss, which has been shown to correlate with the generalization performance of fine-tuned models (ju2022robust). ju2023generalization extend this Hessian measure to graph neural networks, which can guide the design of optimization algorithms to regularize the Hessian of neural networks (ju2023noise).
@@
Multitask Learning Optimization Algorithms. Multitask learning can be viewed as a multiobjective optimization problem (papadimitriou2000approximability), where the goal is to identify the Pareto frontier among multiple objectives (sener2018multi). One common MTL optimization algorithm is to reweight task losses and optimize a weighted combination of task losses (liu2022auto; royer2024scalarization). Our goal is to maximize the averaged prediction performance of all tasks. Thus, we are interested in partitioning the tasks into similar groups, where tasks are closely related within each group and can differ significantly across groups. Another interesting line of work is designing branching neural networks such as tree structures (vandenhende2019branched; guo2020learning), where each layer contains multiple modules to handle different tasks (lu2017fully). Compared with branching methods, task grouping may be more suitable for handling many tasks (like hundreds to thousands). In this regime, negative interference between tasks is almost unavoidable, and clustering tasks into similar groups could provide a more efficient strategy than designing a single neural network that handles all tasks.
@@
Influence Functions. There is a line of work estimating the influence of adding or removing one sample on the whole dataset. Influence functions (koh2017understanding) based on efficient approximation of the Hessian inverse provide one way to approximate this. Random sampling-based approaches to measuring leave-one-out influence have also been studied (ilyas2022datamodels; park2023trak). The distinction between these works and us is we focus on task-level affinity, whereas this literature focuses on estimating the influence of a single data sample.
@@ 2. Preliminaries
Suppose we are interested in making predictions on $n$ tasks. We are given a set of samples for training and testing each task. We aim to design a prediction algorithm to maximize the averaged testing performance over all the $n$ tasks simultaneously. We assume that the samples from all the tasks are supported on a joint product between a $p$-dimensional feature space $\mathcal{X}$ and a label space $\mathcal{Y}$. To precisely discuss task relationships, we formally define what we mean by a multitask learning algorithm.
@@
For any subset $S\subseteq\{1,2,\dots,n\}$, a multitask learning algorithm $f$ takes the training data of all the tasks in $S$ and combines them in a joint training procedure. Then, the (jointly trained) model is tested on each task $t\in S$. In the end, a test result is obtained for each $t$. Let us denote the test result as $f(S,t)$. Thus, the algorithm’s output will include $\left\lvert S\right\rvert$ results for any subset $S$, one for each $t\in S$.
Given a multitask learning algorithm, the transfer between the $n$ tasks can be viewed through the results of $f$, applied to combinations of tasks as subsets. This notion of transfer underlies many existing multitask learning systems. We give two examples below, which are used in prior works to tackle task transfer in complex visual systems (zamir2018taskonomy; standley2020tasks).
@@
Next, we discuss higher-order task affinity, analogous to sampling features in random forests. First, fix an integer $m$, which is the number of subsets we would like to sample (e.g., analogous to the number of decision trees in a random forest). We independently sample $m$ subsets out of the set $\{1,2,\ldots,n\}$, each subset having a size of $\alpha$, chosen uniformly over all such subsets. Let us denoted the $m$ subsets as $S_{1},S_{2},\ldots,S_{m}$. Then, compute $f(S_{k},j)$, for every $k=1,2,\dots,m$, and $j=1,\dots,\alpha$. Lastly, compute $T_{i,j}$ as the average value of $f$ among all subsets including tasks $i,j$:
@@
where $n_{i,j}$ is the number of subsets that include both $i,j$. This leads to another task affinity matrix $T$, which better captures the higher-order relationship among tasks.
@@ 3. Task Affinity Estimation
We now describe a new method for estimating task affinity scores. To circumvent the cost of full-model training, we describe an empirical observation regarding pre-training and fine-tuning. Then, we present our approach to estimating fine-tuned model parameters for task subsets. Additionally, we use random projection to reduce the dimension of the gradients. We provide an error analysis to justify the design of our algorithm.
@@ 3.1. Linearization of Fine-tuned Models
Our method is motivated by the fact that once we pre-train all the $n$ tasks to obtain a meta-initialization, this initialization can provide representations quickly adapted to the remaining tasks. This is based on the premise that the underlying tasks share structural similarities in multitask learning. As the model fine-tuned to a subset of tasks stays in the affinity of the initiation, the fine-tuning procedure behaves like linear models locally.
To illustrate this observation, we consider three scenarios involving graph neural networks (GNNs) and transformers (BERT and T5). We test GNNs on a multi-label prediction dataset on a YouTube graph (yang2012defining), using a 3-layer SIGN network (frasca2020sign). This dataset includes $n=100$ subtasks, one corresponding to the node labels of a subgraph of the whole graph. For transformers, we take a pretrained BERT model and fine-tune it on a sentence classification dataset (yu2020fine), which contains $n=26$ tasks. We also use a pretrained T5-Base model and fine-tune it on a sentence classification dataset with $100$ instructions (bach2022promptsource), which has $n=100$ tasks. In each experiment, we first train a meta-initialization $\theta^{\star}$ by training on all tasks combined. Then, we fine-tune $\theta^{\star}$ on a random subset of the tasks.